OpenStudy (anonymous):

FAN & MEDAL!!! Solve by substitution. y = negative-2 4x − 3y = 18 (0, negative-2) (3, negative-2) (negative-2, negative-2)

OpenStudy (anonymous):

@J-bird

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

k, igreen isn't

OpenStudy (anonymous):

good

OpenStudy (anonymous):

I have a couple questions

OpenStudy (anonymous):

its 3,negative 2

OpenStudy (anonymous):

Which of the following is the solution of the given system of equations? -negativex + y − z = 14 2y + z = negative-12 y = 2 (4, 2, 16) (4, 2, negative-16) (14, 2, negative-16) (negative-4, 2, 21)

OpenStudy (anonymous):

what grade are you in BTW

OpenStudy (anonymous):

11th, it's pathetic I don't know algebra

OpenStudy (anonymous):

i did algebra last year and im in 9th

OpenStudy (anonymous):

you're smart, what's the answer Which of the following is the solution of the given system of equations? -negativex + y − z = 14 2y + z = negative-12 y = 2 (4, 2, 16) (4, 2, negative-16) (14, 2, negative-16) (negative-4, 2, 21)

OpenStudy (anonymous):

why are there 3 digits for answers?? i only see z and y

OpenStudy (anonymous):

oh nvm i see it

OpenStudy (anonymous):

cause there's a -x

OpenStudy (anonymous):

ok so z=8 y=2 with those number can you figure out x??

OpenStudy (anonymous):

and stuff

OpenStudy (anonymous):

get out of here reaves im trying to help him

OpenStudy (anonymous):

hmm hang on a sec

OpenStudy (anonymous):

not really, I don't really have time , I have like 4 more Q's and only 7 mins

OpenStudy (anonymous):

z= -16

OpenStudy (anonymous):

Solve by back substitution. x + 2y + z = 4 x + z = 4 z = 2 (3, 1, negative-2) (2, 0, 2) (negative-1, 0, negative-2) (2, 1, 2)

OpenStudy (anonymous):

z and x both equal 2

OpenStudy (anonymous):

let me find y a sec

OpenStudy (anonymous):

its 2 0 2

OpenStudy (anonymous):

Solve by back substitution. y = negative-2 x + y = negative-4 x + y + 5 = z (0, negative-2, 0) (negative-2, negative-2, 1) (2, negative-2, 0) (1, negative-2, 1)

OpenStudy (anonymous):

-2 -2 0

OpenStudy (anonymous):

CAN YOU SHOW WORK FOR THIS Solve the system by substitution. negative-3 x − 3 y = 3 y = negative-5 x − 17

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

sorry taking so long this one is harder XD

OpenStudy (anonymous):

-3x and -3y. to pass off showing your work, just substitute them into the equation, thats what i did and aced algebra

OpenStudy (anonymous):

If you get the solution 12 = 5 there is no solution to the system of equations. True False

OpenStudy (anonymous):

@J-bird

OpenStudy (anonymous):

true

OpenStudy (anonymous):

last one If you get the solution 1 = 9, there are infinitely many solution to the system of equations. True False

OpenStudy (anonymous):

false

OpenStudy (anonymous):

0=0 is infinitely many solutions

OpenStudy (anonymous):

thanks